OpenStudy (anonymous):

1/x -1/3x=1/2x+1/6x solve equation.

OpenStudy (anonymous):

1/x is clear. 1/3x? Do you mean x/3 (a third of x)? Or 1 over 3x?

OpenStudy (anonymous):

2/3x=4/6x 2/3x=2/3x 1=1

OpenStudy (anonymous):

Hmm... not sure about that, hakkunamtata.

OpenStudy (anonymous):

no? what do you think oh and it is 1/3x its solving ration equations. i think im supposed to get a lcd

OpenStudy (anonymous):

chels10, Is the following the equation?\[\frac{1}{x}-\frac{1}{3 x}=\frac{1}{2 x}+\frac{1}{6 x} \]

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

If you evaluate each side of the equation you get this:\[\frac{1}{x}-\frac{1}{3 x}=\frac{1}{2 x}+\frac{1}{6 x}\to \frac{2}{3 x}=\frac{2}{3 x}\]2/(3x)=2/(3x) is an identity equation. It is true for at least every real number.